# Versatile Demonstration Interface: Toward More Flexible Robot Demonstration Collection

### v1

@@ Abstract
Previous methods for Learning from Demonstration leverage several approaches for a human to teach motions to a robot, including teleoperation, kinesthetic teaching, and natural demonstrations. However, little previous work has explored more general interfaces that allow for multiple demonstration types. Given the varied preferences of human demonstrators and task characteristics, a flexible tool that enables multiple demonstration types could be crucial for broader robot skill training. In this work, we propose Versatile Demonstration Interface (VDI), an attachment for collaborative robots that simplifies the collection of three common types of demonstrations. Designed for flexible deployment in industrial settings, our tool requires no additional instrumentation of the environment. Our prototype interface captures human demonstrations through a combination of vision, force sensing, and state tracking (e.g., through the robot proprioception or AprilTag tracking). Through a user study where we deployed our prototype VDI at a local manufacturing innovation center with manufacturing experts, we demonstrated the efficacy of our prototype in representative industrial tasks. Interactions from our study exposed a range of industrial use cases for VDI, clear relationships between demonstration preferences and task criteria, and insights for future tool design.
@@ I Introduction
When teaching robots new skills, it can be difficult to effectively transfer expert knowledge to a robot’s behavior. Learning from Demonstration (LfD) [1] has emerged as a prevalent strategy for humans to teach such skills, where a person demonstrates a desired motion rather than relying on traditional low-level robot programming. Within LfD, several approaches have emerged for humans to provide demonstrations; including teleoperation, kinesthetic teaching, and natural interfaces [2]. Each of these methods has relative advantages and disadvantages that depend on a number of criteria, including task characteristics, individual demonstrator preferences, and the algorithm that encodes the robot skill. As an example, teleoperation is a convenient approach to collect demonstration data from a distance, but can suffer in terms of demonstration speed and quality depending on the teleoperation input device. On the other hand, natural demonstrations, such as video recordings of humans or instrumented tools, allow humans to show the robot exactly how they would perform the task. However, natural interfaces often suffer from correspondence issues between the human demonstrations and the robot task strategy (e.g., due to configuration limits). Given the varied characteristics and benefits of different demonstration modalities, we believe that future flexible automation systems [3] should make it easier to provide different types of demonstrations. In this work, we propose a demonstration interface (shown in Figure 1) that enables experts to provide demonstrations through three common demonstration modalities.
@@
With recent advancements in robot imitation learning, the need for new tools to collect demonstration data is pertinent. However, there has been little emphasis on creating general-purpose tools for collecting multiple forms of demonstrations. Ideally such general-purpose demonstration tools would be compatible with existing robot and task setups. Creating a single interface that combines demonstration modalities poses a number of fundamental challenges. This work addresses two key issues: how to design the interface for flexible use between demonstration modalities, and how to design the corresponding interaction modalities for data collection and to reduce the effort associated with switching between demonstration types.
This paper presents Versatile Demonstration Interface (VDI), a robot-mounted interface that enables users to provide three types of demonstrations during LfD. The primary contributions of this paper are the development of a prototype VDI interface and a user study with manufacturing experts that assesses its potential in manufacturing tasks. Our results suggest varied use cases for VDI and the corresponding interaction modalities. In the remaining sections, we review related work in the area of LfD, describe the development and features of our prototype, and present the findings from a user study where manufacturing experts performed a set of tabletop tasks using our prototype tool.
@@ II Related Work
Our work develops a flexible LfD interface that combines multiple demonstration modalities. To contextualize our contributions, we provide a brief review of LfD methods and previous methods that assess and combine multiple demonstration modalities.
@@
Learning from Demonstration Methods – From its introduction nearly thirty years ago [4], Learning from Demonstration (also often called programming by demonstration) has gained popularity as a paradigm to teach robots new skills, circumventing many challenges associated with traditional robot programming [1]. Past methods can generally can be categorized into three primary approaches – kinesthetic teaching, teleoperation, and passive observation (i.e., natural demonstrations) [5]. Kinesthetic teaching, where a robot arm is guided through training motions, is commercially available in many collaborative robot platforms and been used to teach robots a variety of complex skills [6, 7, 8, 9]. These methods are designed to be intuitive for end users and prioritize feasibility of skill transfer by inherently limiting the demonstrations to motions that can be achieved by the robot. The usability of such methods is often correlated with the control method, which frequently leverage gravity compensation, compliant, or impedance control [10, 11]. Teleoperation, where a robot is controlled from a distance, is a longstanding robot control paradigm [12] that can also be used to collect demonstrations [13]. More recently, vision-based learning systems have leveraged teleoperation to crowd source demonstration data that resembles deployment environments (i.e., with no human present) [14]. Natural demonstrations, where a person can perform a demonstration without specific considerations for robot execution, is a logical approach to collect high-quality and realistic human data, but poses many challenges such as correspondence (i.e., how to map human demonstrations to robot capabilities and limits) and state tracking [5]. To track the human, many recent methods leverage instrumented tools [15, 16, 17], augmented reality [18, 19], or learning from observations [20, 21].
@@
With recent advances in imitation learning techniques [22, 23], interest in LfD paradigms that can collect data in an effective way has intensified. New physical interfaces have emerged to improve data collection through existing techniques. For example, the Universal Manipulation Interface (UMI) [24] is a low-cost natural demonstration interface for collecting prehensile demonstrations. Mobile ALOHA [25], HATO [26], and GELLO [27] are teleoperation setups to control bimanual robot arms with intuitive input devices across diverse household tasks.
@@
Modality Comparisons and Fusing Modalities – Given the breadth of LfD approaches, previous work has systematically compared demonstration methods. Early comparisons highlighted the value of kinesthetic teaching over more traditional programming approaches (e.g., teach pendants) and teleoperated data collection [28, 29, 30, 31]. More recently, Maric et al. [32] quantified differences in performance and task load when performing contact-rich tasks using kinesthetic teaching and natural demonstrations and Praveena et al. [33] compared teleoperation, kinesthetic, hand, and instrumented tool demonstrations across various pick and place tasks. Generally, these previous studies suggest that participants perform better and prefer using more natural input methods (e.g., kinesthetic teaching over traditional programming, hand tracking input over kinesthetic teaching). However, in some cases, task characteristics may drive the choice of demonstration method. For example, a very precise task might benefit from teleoperation, where a user’s input can be scaled down (e.g., as is popular in robot-assisted surgery [34]). Alternatively, when operating a heavy tool, it may be beneficial for the operator to provide kinesthetic guidance over natural demonstrations to avoid the physical burden of the tool (e.g., through robot force offloading) [35]. When considering a flexible tool that can be used across many tasks, we thus see merit in a tool that combines demonstration modalities.
@@
While limited work has investigated fusing together demonstration modalities, initial results suggest there may be benefits for both user interactions and policy learning. Open Multi-Purpose Interface (OMI) [36] combines kinesthetic teaching and teleoperation and exemplifies the benefits of a flexible interface in a multi-step home-care task using a mobile collaborative robot. In policy learning, MimicPlay [37] highlights the benefits of combining natural demonstrations with a small amount of teleoperated demonstrations to overcome robot correspondence issues. Similarly, PoCo [38] has shown how information from different demonstration modalities can be combined to learn more general skills.
@@ III Versatile Demonstration Interface
To illustrate the opportunities associated with tools combining multiple demonstration modalities and to collect design and use case feedback, we developed an initial prototype of VDI. Here we discuss the important features of our physical and interaction design. Example interactions using our prototype interface are shown in the supplemental video.
@@ III-A Design and Sensing
Our initial physical prototype had a few known design limitations that we plan to address in future versions. First, we included only a single camera that is used at times for data collection and at times for state tracking. For future visuomotor policy learning, it would beneficial to have an additional camera such that there is always a tool-centric view that can be used as part of the demonstration data. Our initial prototype also consists of a wired connection to a box containing the power electronics. We envision future iterations of the tool may include on-board electronics, batteries, wireless communication, and docked charging.
@@ III-B Demonstration Modalities
Teleoperation – In our implementation, the 6D teleoperation input to the robot is provided through a SpaceMouse, though our approach is amenable to any remote input device. Because the SpaceMouse does not provide force feedback to the operator, we operate our robot in an admittance control mode and use LEDs and sound to communicate when the level of applied force is approaching the set force limit. The gains (i.e., scaling) of the teleoperation input generally trade off precision with speed. In our implementation, we empirically tuned the gains for precise manipulations (as was required by the study tasks). The operator is able to switch into the teleoperation mode at any point when the tool is attached by pressing on the remote input device. If the operator exceeds the set force limit, the robot exits the teleoperation mode.
@@
$\operatorname*{arg\,min}_{\textbf{p},\boldsymbol{\omega}}$ $\sum\limits_{i}w_{i}\phi_{i}(\textbf{p},\boldsymbol{\omega})$ (1)
@@
where p is the commanded robot position, $\boldsymbol{\omega}$ is the robot orientation (in the form of a rotation vector), $w_{i}$ are the objective weights, $\phi_{i}$ are the objectives, and $g_{i}$ are the $m$ inequality constraints. In our implementation, the objectives select robot poses that (1) balance the distance between the camera and the tool, (2) a camera angle that centers the tool within the camera view and tracks the AprilTags from a specified angle, and (3) a small term that encourages a neutral (i.e., centered) camera pose and angle that we found empirically (through pilots) helps to avoid local minima in the optimization. The inequality constraints are used to limit the rotation and commanded position of the robot to within the robot’s dexterous workspace. As part of the optimization, we also limited the linear and angular velocity of the commanded robot pose for participant safety. The specific values and terms are in our open-source code [^1] https://github.com/mhagenow01/vdi. The state of the tool is estimated using the AprilTags. Each time a camera sees an AprilTag, the known configuration provides an estimate of the 6D tool pose. While we were able to design our prototype to include explicit state tracking, we are interested in the future to investigate other methods (or ensembles of approaches) for estimating the tool pose (e.g., FoundationPose [47], InfraredTags [48]). The various estimates of the tool pose are used as input to an Extended Kalman Filter (EKF) which provides the final pose estimation. In our implementation, the pose estimation was throttled to 5 Hz to allow for more observations per output pose estimate. For the purposes of our initial prototype implementation and optimizing the camera pose, we found this rate to be sufficient. However, for demonstration data collection, a higher estimation rate will be necessary. We imagine achieving a higher frequency through existing sensor fusion methods (e.g., adding an IMU to the interface). If the robot camera is unable to see any of the tool markers, the system beeps to alert the operator.
@@ IV Evaluation
The primary purpose of our evaluation was to assess the feasibility of a tool that combines demonstration modalities and to assess how it might be applied in real-world manufacturing contexts. Thus, we aimed to collect usability data, design feedback, and to better understand potential use cases. The prototype was not designed for high-fidelity data collection, and thus, our experiment does not focus on data collection nor demonstrating the value of heterogeneous demonstration data in policy learning. We plan to assess these aspects in future studies with an updated interface.
@@ IV-A Task Description
We designed two tasks to resemble processes in industrial manufacturing. Both task setups are shown in Figure 5. The first task was a rolling task, which proxies a manual thermomolding task and consisted of rolling a block of play-doh against a contoured mold. The play-doh was initially formed into a puck-like shape and secured to the mold using a pin (for one-handed completion). The purpose of the first task was to understand the perceived utility of the three demonstration modes in a complex, contact-rich task.
The second task was a press-fitting task where patterns of cylindrical fittings were pressed into a mold. The task was designed to combine coarse, fine, and force-controlled motions. Participants pressed sixteen fittings into a mold. The fittings were arranged in groups of eight with the two groups on opposite ends of the workbench. The purpose of the second study task was to further probe preferences for the demonstration modalities on a task with different characteristics (e.g., less path following, more coarse motion).
IV-B Evaluation with Manufacturing Experts
Our study involved nine experts (7M, 2F), aged 33–69 ($M=45.3,SD=11.1$), recruited from the Berkshire Innovation Center (BIC) via email. A summary of the participants is provided in Table I. Eight of the participants were from a manufacturing background with an average of 15 years of experience. The procedure was administered under a protocol #2402001226 approved by the Institutional Review Board (IRB) of the Massachusetts Institute of Technology. The experiment was designed to last an hour and a small remuneration was paid for participation.

### v2

@@ Abstract
Previous methods for Learning from Demonstration leverage several approaches for a human to teach motions to a robot, including teleoperation, kinesthetic teaching, and natural demonstrations. However, little previous work has explored more general interfaces that allow for multiple demonstration types. Given the varied preferences of human demonstrators and task characteristics, a flexible tool that enables multiple demonstration types could be crucial for broader robot skill training. In this work, we propose Versatile Demonstration Interface (VDI), an attachment for collaborative robots that simplifies the collection of three common types of demonstrations. Designed for flexible deployment in industrial settings, our tool requires no additional instrumentation of the environment. Our prototype interface captures human demonstrations through a combination of vision, force sensing, and state tracking (e.g., through the robot proprioception or AprilTag tracking). Through a user study where we deployed our prototype VDI at a local manufacturing innovation center with manufacturing experts, we demonstrated VDI in representative industrial tasks. Interactions from our study highlight the practical value of VDI’s varied demonstration types, expose a range of industrial use cases for VDI, and provide insights for future tool design.
@@ I Introduction
When teaching robots new skills, it can be difficult to effectively transfer expert knowledge to a robot’s behavior. Learning from Demonstration (LfD) [1] has emerged as a prevalent strategy for humans to teach such skills, where a person demonstrates a desired motion rather than relying on traditional low-level robot programming. Within LfD, several approaches have emerged for humans to provide demonstrations; including teleoperation, kinesthetic teaching, and natural interfaces [2]. Each of these methods has relative advantages and disadvantages that depend on a number of criteria, including task characteristics, individual demonstrator preferences, and sometimes the algorithm that encodes the robot skill. As an example, teleoperation is a convenient approach to collect demonstration data from a distance, but can suffer in terms of demonstration speed and quality depending on the teleoperation input device. On the other hand, natural demonstrations, such as video recordings of humans or instrumented tools, allow humans to show the robot exactly how they would perform the task. However, natural interfaces often suffer from correspondence issues in mapping human demonstrations to the robot’s capabilities. Given the varied benefits of different demonstration modalities, we believe that future flexible automation systems [3] should make it easier to provide different types of demonstrations. In this work, we propose a demonstration interface (shown in Figure 1) that enables experts to provide demonstrations through three common demonstration modalities.
@@
With recent advancements in robot imitation learning, the need for new tools to collect demonstration data is pertinent. However, there has been little emphasis on creating general-purpose tools for collecting multiple forms of demonstrations. Ideally such general-purpose demonstration tools would be compatible with existing robots, task setups, and provide value to experts in important domains (e.g, manufacturing). Creating a single interface that combines demonstration modalities poses a number of fundamental challenges. This work addresses two key issues: how to design the interface for flexible use between demonstration modalities and how to reduce the effort associated with switching between demonstration types.
This paper presents Versatile Demonstration Interface (VDI), a robot-mounted interface that enables users to provide three types of demonstrations during LfD. The primary contributions of this paper are the development of a tangible instantiation of VDI and a user study with manufacturing experts that analyzes the potential of VDI’s interaction modalities in representative manufacturing tasks. Our results suggest that our implementation of VDI was sufficient to demonstrate representative manufacturing tasks and highlights use cases for VDI’s interaction modalities in manufacturing tasks. In the remaining sections, we review related work in the area of LfD, describe the development and features of our interface, and present the findings from a user study where manufacturing experts performed a set of tabletop tasks using the VDI tool.
@@ II Related Work
Our flexible LfD interface combines multiple demonstration modalities. To contextualize our contributions, we provide a brief review of LfD methods and previous methods that assess and combine multiple demonstration modalities.
Learning from Demonstration Methods – From its introduction nearly thirty years ago [4], Learning from Demonstration (also often called programming by demonstration) has gained popularity as a paradigm to teach robots new skills, circumventing many challenges associated with traditional robot programming [1].
@@
Past methods can generally can be categorized into three primary approaches – kinesthetic teaching, teleoperation, and passive observation (i.e., natural demonstrations) [5]. Kinesthetic teaching, where a robot arm is guided through training motions, is commercially available in many collaborative robot platforms and been used to teach robots a variety of complex skills [6, 7, 8, 9]. These methods are designed to be intuitive for end users and prioritize feasibility of skill transfer by inherently limiting the demonstrations to motions that can be achieved by the robot. The usability of such methods is often related to the control method, which frequently leverage gravity compensation, compliant, or impedance control [10, 11]. Teleoperation, where a robot is controlled from a distance, is a longstanding robot control paradigm [12] that can also be used to collect demonstrations [13]. More recently, vision-based learning systems have leveraged teleoperation to crowd source demonstration data that resembles deployment environments (i.e., with no human occulsions) [14]. Natural demonstrations, where a person can perform a demonstration without specific considerations for robot execution, is a logical approach to collect high-quality and realistic human data, but poses many challenges such as correspondence (i.e., how to map human demonstrations to robot capabilities and limits) and state tracking [5]. To track the human, many recent methods leverage instrumented tools [15, 16, 17], augmented reality [18, 19], or learning from observations [20, 21].
@@
With recent advances in imitation learning [22, 23], interest in LfD paradigms for effective data collection have intensified. Recent efforts include new physical interfaces to improve data collection. For example, the Universal Manipulation Interface (UMI) [24] is a low-cost natural demonstration interface for collecting prehensile demonstrations. Mobile ALOHA [25], HATO [26], and GELLO [27] are teleoperation setups to control bimanual robot arms with intuitive input devices across diverse household tasks.
@@
Modality Comparisons and Fusing Modalities – Given the breadth of LfD approaches, previous work has systematically compared demonstration methods. Early comparisons highlighted the value of kinesthetic teaching over more traditional programming approaches (e.g., teach pendants) and teleoperated data collection [28, 29, 30, 31]. More recently, Maric et al. [32] quantified differences in performance and task load when performing contact-rich tasks using kinesthetic teaching and natural demonstrations and Praveena et al. [33] compared teleoperation, kinesthetic, hand, and instrumented tool demonstrations across various pick and place tasks. Generally, these previous studies suggest that participants perform better and prefer using more natural input methods (e.g., kinesthetic teaching over traditional programming, hand tracking input over kinesthetic teaching). However, in some cases, task characteristics may drive the choice of demonstration method. For example, a very precise task might benefit from teleoperation, where a user’s input can be scaled down (e.g., as is popular in robot-assisted surgery [34]). Alternatively, when operating a heavy tool, it may be beneficial for the operator to provide kinesthetic guidance over natural demonstrations to avoid the physical burden of the tool (e.g., through robot force offloading) [35]. When considering a flexible tool for use in many tasks, we thus see merit in interfaces combining modalities.
@@
While limited work has investigated tools fusing together demonstration modalities, initial results suggest there may be benefits for both user interactions and policy learning. Open Multi-Purpose Interface (OMI) [36] combines kinesthetic teaching and teleoperation and exemplifies the benefits of a flexible interface in a multi-step home-care task using a mobile collaborative robot. In policy learning, MimicPlay [37] highlights the benefits of combining natural demonstrations with a small amount of teleoperated demonstrations to overcome robot correspondence issues. Similarly, PoCo [38] has shown how information from different demonstration modalities can be combined to learn more general skills. While heterogeneous learning is not a focus of this work (we instead focus on human perceptions and use cases for multiple forms of demonstration), we are optimistic that advances in heterogeneous learning will open further opportunities for more general-purpose tools for human demonstrations.
@@ III Versatile Demonstration Interface
To illustrate the opportunities associated with combining multiple demonstration modalities and to collect design and use case feedback, we developed a hardware implementation of VDI. Here we discuss the important features of our physical and interaction design. Example interactions using our prototype interface are shown in the supplemental video.
@@ III-B Demonstration Modalities
Teleoperation – In our implementation, the 6D teleoperation input to the robot is provided through a SpaceMouse, though our approach is amenable to any remote input device. Because the SpaceMouse does not provide force feedback to the operator, we operate our robot in an admittance control mode and use LEDs and sound to communicate when the level of applied force is approaching the set force limit. The gains (i.e., scaling) of the teleoperation input generally trade off precision with speed. In our implementation, we empirically tuned the gains for precise manipulations (as was required by the study tasks). The operator can enter teleoperation mode at any point when the tool is attached by pressing on the remote input device. If the operator exceeds the set force limit, the robot exits the teleoperation mode.
@@
$\operatorname*{arg\,min}_{\textbf{p}_{r},\textbf{R}_{r}}$ $\sum\limits_{i}w_{i}\phi_{i}(\textbf{p}_{r},\textbf{R}_{r},\textbf{p}_{h},\textbf{R}_{h})$ (1)
@@
where $\textbf{p}_{r}$ is the commanded position of the robot camera, $\textbf{R}_{r}$ is the robot camera rotation matrix, $\textbf{p}_{h}$ is the estimated position of the human interface, $\textbf{R}_{h}$ is the estimated rotation matrix of the human interface, $w_{i}$ are the objective weights, $\phi_{i}$ are the objectives, and $g_{i}$ are the $m$ inequality constraints. While many viewpoint heuristics are possible, we leveraged four objectives that yielded safe (for the user study) and appropriate behaviors. We provide the objectives and parameters for reproducibility, but acknowledge that different robots and workspaces may require different objectives or parameters. Our robot camera pose objectives optimize:
  • Maintaining an appropriate distance between the robot camera and the human interface.
    $\phi_{1}(\textbf{p}_{r},\textbf{R}_{r},\textbf{p}_{h},\textbf{R}_{h})=\left((\textbf{R}_{r}(\textbf{p}_{h}-\textbf{p}_{r}))^{\textrm{T}}\hat{\textbf{z}}-d\right)^{2}$ (2)
    where $\hat{\textbf{z}}=[0,0,1]^{\textrm{T}}$ is a unit vector that extracts the distance relative to the camera z-axis and $d$ is the desired camera-interface distance. In our implementation, $w_{1}=100$ and $d=0.3\textrm{m}$. The distance was selected to balance field-of-view and AprilTag recognition.
  • Centering the tool in the camera view by aligning the camera with the vector between the camera and interface.
    $\phi_{2}(\textbf{p}_{r},\textbf{R}_{r},\textbf{p}_{h},\textbf{R}_{h})=\left((\frac{\textbf{p}_{h}-\textbf{p}_{r}}{||\textbf{p}_{h}-\textbf{p}_{r}||})^{\textrm{T}}(\textbf{R}_{r}\hat{\textbf{z}})-1\right)^{2}$ (3)
    In this objective, $w_{2}\!=\!100$. We also considered an additional objective that constrained the viewing angle of the centered tool (for better viewing of the AprilTags), but found it had little impact on behavior empirically.
  • Encouraging a neutral (i.e., centered) camera pose and angle that we found through pilots helps to avoid local minima in the optimization.
    $\phi_{3}(\textbf{p}_{r},\textbf{R}_{r},\textbf{p}_{h},\textbf{R}_{h})=||\textbf{p}_{r}-\textbf{p}_{n}||^{2}$ (4)
    $\phi_{4}(\textbf{p}_{r},\textbf{R}_{r},\textbf{p}_{h},\textbf{R}_{h})=\angle(\textbf{R}_{n}^{\textrm{T}}\textbf{R}_{r})^{2}$ (5)
    where $\textbf{p}_{n}=[0,-0.4,0.35]^{\textrm{T}}$ is the neutral camera position, $\textbf{R}_{n}$ is the neutral orientation, and ($\angle$($\cdot$)∠ ( ⋅ ) the magnitude of rotation. In our experimental setup, we only found it empirically beneficial to penalize rotation about the $y$-axis, which corresponded to the principal axis of the robot’s workspace (i.e., table). We encouraged the neutral camera pose through small penalties ($w_{3}\!=\!2$ and $w_{4}\!=\!0.5$).
[…]
@@ IV Evaluation
Our vision is to enable flexible Learning from Demonstration interfaces that can adapt to different demonstration needs (by enabling multiple demonstration modalities) and do not require sensing beyond what is available onboard the robot and LfD interface. To assess our prototype robot interface, we conducted a user study with manufacturing experts. The goals of our study were to (1) assess expert interactions with the robot-tracking based natural demonstration mode to understand its utility and limitations, (2) to probe participants regarding manufacturing applications where each demonstration mode would be valuable (i.e., demonstrating the practical value of a single flexible interface for multiple demonstrations in manufacturing) and (3) to further understand opportunities for manufacturing tasks that would benefit from VDI’s ability to easily switch between demonstration modes within a task.
@@ IV-A Task Description
We designed two tasks to resemble processes in industrial manufacturing. Both task setups are shown in Figure 4. The first task was a rolling task, which proxies a manual thermomolding task and consisted of rolling a block of play-doh against a contoured mold. The play-doh was initially formed into a puck-like shape and secured to the mold using a pin (for one-handed completion). The purpose of the first task was to expose participants to using the three demonstration modes of VDI in a complex, contact-rich task.
The second task was a press-fitting task where patterns of cylindrical fittings were pressed into a mold. The task was designed to combine coarse, fine, and force-controlled motions. Participants pressed sixteen fittings into a mold. The fittings were arranged in groups of eight with the two groups on opposite ends of the workbench. The purpose of the second study task was for participants to explore switching between demonstration modalities and to use VDI under varied task characteristics (e.g., coarser, discrete motions).
IV-B Manufacturing Study Population
Our study involved nine experts, summarized in Table I, (7M, 2F), aged 33–69 ($M=45.3,SD=11.1$), recruited from the Berkshire Innovation Center via email. Eight participants were from a formal manufacturing background with an average of 15 years of experience. The procedure was administered under a protocol #2402001226 approved by the Institutional Review Board (IRB) of the Massachusetts Institute of Technology. The experiment was designed to last an hour and a small participation remuneration was paid.